A stone initially moving at 8.0 m/s on a level surface comes to rest due to friction after it travels 11 m. What is the coefficient of kinetic friction between the stone and the surface?

answer: 0.30

Please show work
Answer #1

acceleration = (v^2 -vo^2)/2d = [0- (8.0 m/s)^2 ]/ 2(11 m) = -2.9 m/s^2 (negative acceleration since the object is slowing down).

To find the coefficient of friction we'll use Newton's second law of motion: F = ma. Here the applied force is the force of friction. Therefore (mu) (m)(g) = ma. As you see mass "m" can be cancelled and since we know the acceleration we can easily calculate the coefficient of friction. Answer: (mu) = 0.29